OpenStudy (anonymous):

factor 4-a^2-2ab-b^2

OpenStudy (anonymous):

4-a^2-2ab-b^2 =4 - (a^2+2ab+b^2) = (2)^2 - (a+b)^2 =(2-a-b)(2+a+b) =-(a+b-2)(a+b+2)